# Training Gradient Boosted Decision Trees on Tabular Data Containing Label Noise for Classification Tasks

### v1

@@ Abstract
Label noise refers to the phenomenon where instances in a data set are assigned to the wrong label. Label noise is harmful to classifier performance, increases model complexity and impairs feature selection. Addressing label noise is crucial, yet current research primarily focuses on image and text data using deep neural networks. This leaves a gap in the study of tabular data and gradient-boosted decision trees (GBDTs), the leading algorithm for tabular data. Different methods have already been developed which either try to filter label noise, model label noise while simultaneously training a classifier or use learning algorithms which remain effective even if label noise is present. This study aims to further investigate the effects of label noise on gradient-boosted decision trees and methods to mitigate those effects. Through comprehensive experiments and analysis, the implemented methods demonstrate state-of-the-art noise detection performance on the Adult dataset and achieve the highest classification precision and recall on the Adult and Breast Cancer datasets, respectively. In summary, this paper enhances the understanding of the impact of label noise on GBDTs and lays the groundwork for future research in noise detection and correction methods.
@@ 1 Introduction
In the area of machine learning and data-driven decision-making, accurate predictive modeling is crucial for diverse applications, from healthcare diagnostics to financial risk assessment. Among various techniques, gradient boosted decision trees (GBDTs) are the state-of-the-art algorithm for tabular data, consistently outperforming deep learning approaches (Grinsztajn et al., 2022).
However, real-world application of GBDTs poses several challenges. One particularly significant challenge is the presence of label noise in the data. Label noise refers to the phenomenon where instances in a dataset are assigned the wrong label. This differs from feature noise, which refers to inaccuracies in the observed feature values of an instance, rather than errors in the assigned labels (Zhu and Wu, 2004). While both feature and label noise can impair model performance, label noise is especially detrimental due to its strong impact on learning, given that there is only one label in contrast to multiple features (Frénay and Verleysen, 2014). Reliable labels are often costly and time intensive to acquire (Frénay and Verleysen, 2014). Although label noise detection is closely related to outlier and anomaly detection, it’s important to note that a mislabeled instance isn’t necessarily an outlier or anomaly and vice versa.
Label noise naturally occurs when human experts are involved (Frénay and Kaban, 2014). The provided information may be insufficient to perform reliable labeling, the expert may simply have made a mistake or the labeling task may leave room for interpretation, so that there is variability in the labeling between experts. Sometimes, communication or encoding problems can also be the cause of incorrect labels.
Feature noise as well as label noise deteriorates classification performance. Additionally, label noise can increase model complexity, increase the number of necessary training instances and complicate the identification of relevant features. Due to these problems, model comparison can be biased if the test set is polluted as well (Frénay and Verleysen, 2014).
1.1 Research Objective
Even though GBDTs outperform deep neural networks (DNNs) on tabular data (Grinsztajn et al., 2022), there are few studies exploring the effects of label noise on GBDTs and means of correcting for label noise (Ponti et al., 2022; Xiang et al., 2020; Brophy et al., 2023). Most of the new label noise research instead focuses on the application of DNNs on image or text data (Li et al., 2020; Jindal et al., 2017), despite the fact that tabular data is the most common data type (Shwartz-Ziv and Armon, 2022).
Therefore, the aim of this work is to
@@ 1.1 Research Objective
    provide an overview of state-of-the-art methods for detecting and handling label noise
@@
    improve upon current state-of-the-art methods by implementing label-noise robust GBDT model
@@
    evaluate the performance of the label-noise robust GBDT model by comparing it to state-of-the-art classifiers
1.2 Structure
To offer an overview of the state of label noise research and to set the stage for this work’s contribution, this section is followed by a review of related work (Section 2). Next, the “Preliminaries” section (Section 3) provides a formulation of the problem of training a supervised machine learning model on data containing label noise, the theoretical background for training GBDTs as well as a taxonomy of label noise. After that, the “Methodology” section (Section 4) outlines the implemented noise detection methods, how they are combined with noise correction methods and details the datasets and evaluation metrics used in this work. The “Experiments” section (Section 5) describes the experimental setup including how label noise is introduced into the data and which programming libraries and model hyperparameters are used. This section initially explores the impact of label noise on GBDTs and subsequently evaluates the performance of various noise detection and correction methods. The results from the experiments are then analyzed and discussed in the “Results and Discussion” section (Section 6). Finally, the “Conclusion” section (Section 7) synthesizes the findings, summarizes this work’s essential contributions and highlights avenues for future research.
@@ 2 Related work
This section presents research related to label noise and first covers the three main types of approaches when dealing with label noise: label noise robust models, label noise tolerant learning algorithms and data cleansing methods (Frénay and Verleysen, 2014). One paragraph is dedicated to the current landscape of label noise research, which involves training DNNs on text or image data. The strengths and weaknesses of each approach are also discussed. After that, the focus is drawn to research investigating training GBDTs on tabular data containing label noise as well as the fact that this type of research is limited.
The last section details how this work aims to fill that gap and derives the primary research questions of this work. It highlights the importance to understand the effects of label noise on GBDTs, to assess the applicability of deep neural network label noise detection methodologies on GBDTs, and to gauge the potential of these adapted GBDTs on benchmark datasets.
2.1 Label Noise Robust Models
Label noise robust models are based on algorithms that aren’t gravely affected by label noise in the first place. These models remain effective even if label noise isn’t cleaned or modeled. Dietterich (2000) showed that bagging generally performs better than boosting when label noise is present. Adaboost is a gradient boosting algorithm that increases the instance weights of misclassified instances (Géron, 2019).
The idea of label noise robust models promises that no or few considerations for label noise need to be made to train an effective model. Those few considerations usually come in the form of label noise robust loss functions and conventional regularization techniques. In practice, however, those models aren’t really that robust to label noise and turned out to only be adequate for simple noise settings and datasets. Also, these methods often have difficulty differentiating mislabeled from hard to learn training samples (Veit et al., 2017).
2.2 Label Noise Tolerant Learning Algorithms
Label noise tolerant learning algorithms either model label noise or regularize models to reduce their sensitivity to noisy instances. Often, the label noise model and classifier are trained simultaneously. Probabilistic approaches assume instances to be generated according to several probability distributions (Frénay and Verleysen, 2014). Bayesian probabilistic approaches use Bayesian priors to model the mislabeling probabilities (Gaba, 1993). But the priors should be chosen carefully, because the results depend on the quality of the chosen prior distribution. Frequentist approaches often use mixture models, where the data is assumed to originate from a mixture of a normal distribution and an “anomalous” distribution (Mansour and Parnas, 1998). Han et al. (2007) sort instances by their prediction scores and assume instances with the most mixed degree of positive and negative instances to be noisy and remove them using a revised EM-algorithm.
Probabilistic and frequentist approaches are grounded in a more theoretical approach than robust or data cleansing methods. But formulating the distribution of noisy and true labels based on specific assumptions may limit the effectiveness with more complex label noise (Li et al., 2019).
2.3 Data Cleansing Methods
Data cleansing methods aim to remove or relabel mislabeled instances to improve the performance of subsequent classifiers. One idea is to implement measures of anomaly (Sun et al., 2007) or measures of the effect on model complexity (Gamberger and Lavrac, 1997) and remove instances that exceed a certain threshold. Jeatrakul et al. (2010) simply remove instances misclassified by a classifier and Verbaeten and Assche (2003) discard instances with too large weights. There are also many heuristics based on k-nearest-neighbors or graphs. For example, removing instances whose label is different from the majority label in its k-neighborhood (Wilson, 1972).
Removing mislabeled instances rather than trying to relabel tends to be more efficient (Cuendet et al., 2007), except when the dataset is imbalanced, because underrepresented classes are more likely to be removed (Frénay and Verleysen, 2014).
Deep learning research often makes use of the so-called “small-loss-trick”, where small-loss samples from the training set are assumed to be clean and selected to update the network robustly (Han et al., 2018; Jiang et al., 2018; Li et al., 2020). Zheng et al. (2020) measure the ratio between a DNN classifier’s confidence on the noisy label and its own label prediction and flip the label to its prediction if the ratio exceeds a certain threshold. Additionally, a theoretical error bound is provided. Pleiss et al. (2020) measure the area under the margin to determine whether an instance is noisy or clean.
Data cleansing methods are often based on simple heuristics which are easy to implement. However, there is a trade-off between relabeling and filtering. Relabeling may introduce more errors, which accumulate in the process of training. Removing might delete minority classes first or instances which are hard to learn but informative, i.e. instances near a decision boundary (Johnson and Khoshgoftaar, 2022).
2.4 Deep Learning and GBDTs
@@
In deep learning, common practices include adding robust architectures that model the noise transition matrix, regularization, improved loss functions and data cleansing, where instances with a likely true label are chosen to update the network (Song et al., 2020). To mitigate the accumulation of errors when further training is based on the network’s own predictions, some researchers have investigated the use of multiple networks (Chen et al., 2023). Co-teaching (Han et al., 2018) involves two neural networks, where each one selects low-loss instances for the other to use in backpropagation. Decouple (Malach and Shalev-Shwartz, 2017) updates the parameters of two networks only when their predictions are in disagreement to avoid overfitting to noisy instances. DivideMix (Li et al., 2020) operates by having each network discard the labels of high-loss instances and then provides the resulting partially-labeled data to the other network for semi-supervised training.
@@
Deep learning achieves state-of-the-art performance on image and text data, but the methods often suffer from slow convergence, high computational cost due to an increased number of hyperparameters or large models due to deep architectures (Song et al., 2020).
Most modern label noise research is applied to image and text datasets using deep neural networks (DNNs) (Li et al., 2020; Jindal et al., 2017) and few on tabular datasets like those mentioned below. Architectures like TabNet (Arik and Pfister, 2021) have been developed to handle tabular data using deep neural networks, which outperform fully connected networks and other networks on several benchmarks, but still don’t achieve state-of-the-art performance as GBDTs do (Borisov et al., 2021).
So far, the research on applying GBDTs on data containing noisy labels is rather sparse, some exceptions include the following: Ponti et al. (2022) calculate the confidence, correctness and variability of the model’s predictions during the GBDT training process to detect and remove mislabeled instances, thereby improving the quality of the dataset. Brophy et al. (2023) adapt three existing methods for estimating the influence of training samples on predictions to GBDTs, which can be used to identify and correct mislabeled or problematic training samples. Xiang et al. (2020) use random label flipping and random feature subspaces to improve label noise robustness and increase the diversity of GBDTs and then aggregate them using bagging.
2.5 Summary and Research Questions
Common practices for dealing with label noise include modeling the noise, increasing the model’s robustness through regularization or robust loss functions or cleaning the data by removing, relabeling or reweighting noisy candidates. DNN-based techniques have received the most attention in recent years and therefore reached state-of-the-art performance on text and image datasets.
Tabular data is the most common data type in real-world applications (Shwartz-Ziv and Armon, 2022), but hasn’t been thoroughly studied in the context of learning with label noise. And tree-based methods like GBDTs remain the best algorithms for learning from tabular data (Grinsztajn et al., 2022). Therefore this paper studies the intersection of these two areas, examining effective strategies for learning from tabular data with label noise using GBDTs, with a focus on data cleansing methods and classification tasks.
Ponti et al.’s (2022) work focuses on data cleansing to combat label noise when training GBDTs. So far only the removal of noisy candidate instances was considered. Therefore this work continues their work by applying automatic relabeling of noisy candidate instances. Zheng et al.’s (2020) and Pleiss et al.’s (2020) data cleansing methods are widely cited and were selected to be adapted to GBDTs. The small-loss-trick (Li et al., 2020) and identifying noisy instances using the weights of boosting algorithms (Verbaeten and Assche, 2003) are also combined into a novel label noise detection method. All the above-mentioned approaches for noisy label detection and correction are implemented to improve the performance of GBDTs on data containing label noise.
Thus, this work aims to answer questions which are concerned with : (1) studying the effects of label noise on conventional GBDTs and (2) comparing the performance of different noise treatment methods implemented for this work. Regarding the first category, the questions are as follows:
  • RQ1. How many epochs of training does it take for noise to negatively impact model performance, i.e. how label noise robust are GBDTs naturally?
  • RQ2. What types of errors does the model make as training progresses?
  • RQ3. Are there certain types of noise which are harder to learn with?
For the second category, the research questions are as follows:
  • RQ4. How well do the detection methods detect noise?
  • RQ5. How do the detection methods change dependent on the noise rate?
  • RQ6. How do the detection metrics change over the course of training?
  • RQ7. Which correction method yields the highest model performance?
@@ 3 Preliminaries
This section provides an overview of the relevant theoretical foundations on which this paper builds upon. Each section focuses on a subject area which is needed to answer the research questions specified above.
The first section explains the basic idea of gradient boosting and how GBDT ensembles apply it by fitting shallow decision trees to the residual errors of their predecessor.
Lastly, the three types of label noise as characterized by Frénay and Verleysen (2014) are detailed, along with their implications for classification models. The types of noise injected into datasets within this work are also mentioned, along with the reasons for their selection. This section serves as a primer, equipping readers with the necessary theoretical knowledge to navigate the subsequent sections.
3.1 Training a GBDT Ensemble
The basic idea of Gradient Boosting is to train predictors sequentially, with each one aiming to fit the residual errors made by its predecessor (Géron, 2019). In other words, the sum of multiple weak learners $m_{t}$, which in the case of GBDTs are shallow decision trees, is supposed to approximate the target $y$ (Brophy et al., 2023; Si et al., 2017):
$y=\sum_{t=1}^{T}m_{t}(x),$ (1)
which can also be expressed as an additive model $f^{t}$ which we stop training once it reaches adequate performance:
$y$ $=f^{t}(x)=f^{t-1}(x)+m_{t}(x),$ (2)
$\Longleftrightarrow m_{t}(x)$ $=y-f^{t-1}(x).$ (3)
This means that $m_{t}$ fits the residual of the previous model $f^{t-1}$. To avoid overfitting to outliers, $m_{t}$ is usually weighted by a parameter called the learning rate $\eta$:
$f^{t}(x)=f^{t-1}(x)+\eta\cdot m_{t}(x).$ (4)
To generalize to any loss function, the negative gradient $g$ of the loss function $l$ with respect to $f^{t-1}$ is instead optimized, and is commonly referred to as the “pseudo-residual”. Using the gradient has the added benefit that it points to the direction of the steepest slope, rather than merely reflecting the magnitude of the difference, as the residual does:
$f^{t}(x)$ $=f^{t-1}(x)+\eta\cdot(-g_{t-1}(x,y)),$ (5)
$g_{t}(x,y)$ $=\frac{\partial l(f^{t}(x),y)}{\partial f^{t}(x)}.$ (6)
If, as an example, the mean squared error is the loss function of choice, then the next weak learner $m_{t}$ would be trained to fit:
$L_{MSE}$ $=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-f^{t}(x_{i}))^{2},$ (7)
$g_{t}(x_{i},y_{i})$ $=\frac{\partial L_{MSE}}{\partial f^{t}(x_{i})}=-\frac{2}{N}(y_{i}-f^{t}(x_{i})),$ (8)
$f^{t}(x_{i})$ $=f^{t-1}(x_{i})+\eta\cdot m_{t}(x_{i}),$ (9)
$=f^{t-1}(x_{i})+\eta\cdot(-g_{t-1}(x_{i},y_{i})),$ (10)
$=f^{t-1}(x_{i})+\eta\cdot\left[-\left(-\frac{2}{N}(y_{i}-f^{t-1}(x_{i}))\right)\right],$ (11)
$=f^{t-1}(x)+\eta\cdot\frac{2}{N}(y_{i}-f^{t-1}(x_{i})),$ (12)
$\Longleftrightarrow m_{t}(x_{i})$ $=\frac{2}{N}(y_{i}-f^{t-1}(x_{i})).$ (13)
This shows that in the case of using the mean squared error as the loss function, the negative gradient is proportional to the residual of the previous weak learner $y_{i}-f^{t-1}(x_{i})$.
3.2 Taxonomy of Label Noise
Label noise can be categorized into three types: Noisy completely at random (NCAR), noisy at random (NAR) and noisy not at random (NNAR) (Frénay and Verleysen, 2014; Song et al., 2020). The corruption process is typically modeled with a noise transition matrix $S\in[0,1]^{c\times c}$, where $S_{ij}:=p(\tilde{y}=j|y=i)$ is the probability of the true label $i$ being flipped into a corrupted label $j$ and $c$ denotes the number of classes. This means that, starting from a noise rate of 50% and up, misclassification is just as or more likely than the true class.
@@ 3.2 Taxonomy of Label Noise
NCAR noise is totally random and depends on neither instance features nor its true class. This type of noise is inherently symmetric, often termed “symmetric noise,” as it equally distributes the probability of the true label being flipped to any other label. With a noise rate $\tau\in[0,1]$, the noise transition matrix $(S_{ij})$ is defined as $S_{ii}=1-\tau$ and $S_{ij}=\frac{\tau}{c-1}$ for $i\neq j$.
NAR noise, also known as class-dependent noise, is independent of instance features but depends on its class, which allows modeling of asymmetric label noise. In that case, the true label is more likely to be mislabeled into a particular label than others, i.e. $S_{ii}=1-\tau$, but $S_{ij}\neq S_{ik}$ for some $k\neq i,j\neq i$. In a stricter case, called pair noise, the true label is flipped only to one certain label. Thus NAR noise can model the case where similar classes are easily confused with one another. For example, a “bird” is more likely to be confused with a “plane” than a “dog”.
NNAR noise, also known as instance- or feature-dependent noise, is deemed to be the most realistic, because it depends on both the instance’s features and its class. This could manifest as instances being more likely to be mislabeled because they are similar to instances of another class or certain regions of the feature space being more likely to be mislabeled. And some classes might be more likely misclassified as with NAR noise. Therefore the corruption probability is defined as $p(\tilde{y}=j|y_{i},x)$ where the data feature $x$ now also affects the chance of the instance being mislabeled.
As an example, Figure 1 shows how the noise transition matrices would look like on a dataset with four classes with no noise, 20% symmetric and 20% pair noise, respectively. In the first noise transition matrix, there is a 0% chance for each class to be switched to another class (excluding the true class). In the second noise transition matrix, there is about a 6.7% chance for each class to be switched to each other class, or a 20% chance in total to be switched to another class. In the last noise transition matrix, there is a 20% chance for each class to be switched to the next class, but a 0% chance to be switched to any other class.
@@
Given the challenges in modeling NNAR noise, as evidenced by the varied approaches to generating it (Zhang et al., 2021; Menon et al., 2018), and to increase the comparability to other works, this work focuses on introducing NCAR and NAR noise to the selected datasets. In the following sections, the noise types will be referred to as “symmetric” and “pair” noise, respectively.

### v2

@@ Abstract
Label noise, which refers to the mislabeling of instances in a dataset, can significantly impair classifier performance, increase model complexity, and affect feature selection. While most research has concentrated on deep neural networks for image and text data, this study explores the impact of label noise on gradient-boosted decision trees (GBDTs), the leading algorithm for tabular data. This research fills a gap by examining the robustness of GBDTs to label noise, focusing on adapting two noise detection methods from deep learning for use with GBDTs and introducing a new detection method called Gradients. Additionally, we extend a method initially designed for GBDTs to incorporate relabeling. By using diverse datasets such as Covertype and Breast Cancer, we systematically introduce varying levels of label noise and evaluate the effectiveness of early stopping and noise detection methods in maintaining model performance. Our noise detection methods achieve state-of-the-art results, with a noise detection accuracy above 99% on the Adult dataset across all noise levels. This work enhances the understanding of label noise in GBDTs and provides a foundation for future research in noise detection and correction methods.
@@ 1 Introduction
Accurate predictive modeling is essential in machine learning applications such as healthcare diagnostics and financial risk assessment. For tabular data, gradient boosted decision trees (GBDTs) are among the top-performing algorithms, often surpassing deep learning models in both accuracy and efficiency (Grinsztajn et al., 2022). However, one major challenge in deploying GBDTs in real-world settings is the presence of label noise—incorrect labels assigned to training instances—which can degrade model performance significantly.
Label noise differs from feature noise, as it directly impacts the learning process by corrupting the target variable rather than the feature set. Mislabeled data, often introduced through human error, is costly to eliminate and typically harder to detect than feature noise (Frénay and Verleysen, 2014). This issue is particularly pressing for GBDTs on tabular data, where research on label noise detection and correction is limited, with most label noise research focusing on training deep learning models on image or text data (Li et al., 2020; Jindal et al., 2016).
This work aims to bridge this gap by exploring label-noise robust methods for GBDTs in tabular data settings. Specifically, we seek to (1) provide a comprehensive overview of state-of-the-art label noise handling techniques, (2) enhance these techniques by developing a robust GBDT model for noisy labels, and (3) evaluate the model’s performance relative to existing classifiers.
Following the introduction, Section 2 provides an overview of label noise research with a focus on deep learning and GBDTs. Section 3 discusses the problem formulation and taxonomy of label noise, while Section 4 explains the methodology, including noise detection and correction techniques. Section 5 describes the experimental setup, covering noise generation, evaluation metrics, and model configuration. Results are presented in Section 6, and Section 7 offers an analysis of the findings. Lastly, Section 8 summarizes the main conclusions and outlines future research directions.
The main contributions are as follows:
@@
    This study addresses the gap in label noise research specifically focused on gradient-boosted decision trees (GBDTs) for tabular data.
@@
    Two noise detection methods designed for deep neural networks were adapted for GBDTs, and a novel detection method, Gradients, was introduced.
@@
    Experiments reveal that GBDTs are naturally robust to label noise, particularly symmetric noise, with early stopping improving performance.
  • 4.
    State-of-the-art noise detection accuracy is achieved with AUM and LRT, both of which can reliably estimate the level of noise in a dataset.
@@ 2 Related Work
This section reviews primary approaches for handling label noise and emphasizes recent developments in deep learning and gradient-boosted decision trees (GBDTs).
2.1 Approaches to Label Noise
Label noise handling typically falls into three main categories: robust models, noise-tolerant learning algorithms, and data cleansing methods (Frénay and Verleysen, 2014). Label noise robust models, such as bagging or variations of Adaboost, are based on algorithms that aren’t gravely affected by label noise in the first place; however, their effectiveness is often limited to simpler noise patterns (Dietterich, 2000; Veit et al., 2017). Noise-tolerant algorithms, including probabilistic and frequentist approaches, either model label noise or regularize models to reduce their sensitivity to noisy instances (Gaba, 1993; Mansour and Parnas, 1998). Lastly, data cleansing methods, which identify and remove or relabel noisy instances, are widely used for their simplicity but risk excluding valuable data or introducing further errors when relabeling (Sun et al., 2007; Han et al., 2018).
2.2 Deep Learning and Label Noise
@@
In deep learning, common practices include adding robust architectures that model the noise transition matrix, regularization, improved loss functions and data cleansing, where instances with a likely true label are chosen to update the network (Song et al., 2020). To mitigate the accumulation of errors when further training is based on the network’s own predictions, some researchers have investigated the use of multiple networks, where each network selects low-loss instances for the other to use in backpropagation (de Vos et al., 2023). Robust DivideMix
(Zhang et al., 2024) operates by having each network discard the labels of high-loss instances and then provides the resulting partially-labeled data to the other network for semi-supervised training. Additionally, NRAT (Chen et al., 2024), a robust adversarial training method, combines a robust loss function and enhanced regularization to improve adversarial robustness under label noise.
@@
Deep learning achieves state-of-the-art performance on image and text data, but the methods often suffer from slow convergence, high computational cost due to an increased number of hyperparameters or large models due to deep architectures (Song et al., 2020). Techniques specifically designed for tabular data remain scarce, as DNN-based methods are typically optimized for unstructured data types.
2.3 GBDTs and Label Noise in Tabular Data
GBDTs are widely regarded as the leading algorithm for tabular data, offering significant efficiency and performance advantages over DNNs in structured datasets (Grinsztajn et al., 2022). Despite this, research on how GBDTs handle label noise remains relatively sparse. Ponti et al. (2022) introduced an approach to detect and remove mislabeled instances in GBDTs by calculating the confidence, correctness, and variability of predictions during training, enhancing dataset quality. Brophy et al. (2023) adapted three established methods for estimating the influence of training samples on predictions, which can help identify correct, mislabeled or problematic samples. Sztukiewicz et al. (2024) examined loss design techniques from deep learning to improve the robustness of decision trees, while Zhu et al. (2024) leveraged isolation forests to detect noisy instances, which were subsequently relabeled using a semi-supervised learning algorithm.
@@ 3 Preliminaries
Label noise is commonly categorized into three types: Noisy Completely at Random (NCAR), Noisy at Random (NAR), and Noisy Not at Random (NNAR) (Frénay and Verleysen, 2014; Song et al., 2020).
@@
In NCAR noise, labels are flipped randomly, independent of instance features or class, leading to symmetric misclassification. This can be modeled with a noise transition matrix $S\in[0,1]^{c\times c}$, where $S_{ij}:=p(\tilde{y}=j|y=i)$ represents the probability of the true label $i$ being misclassified as label $j$, and $c$ denotes the number of classes. For NCAR noise with rate $\tau\in[0,1]$, the transition matrix is defined as $S_{ii}=1-\tau$ for correct labels and $S_{ij}=\frac{\tau}{c-1}$ for $i\neq j$.
NAR noise, also called class-dependent noise, introduces asymmetry in the label flipping process. In this case, the mislabeling probability depends on the class but remains independent of instance features. Thus, we have $S_{ii}=1-\tau$, while $S_{ij}\neq S_{ik}$ for certain $k\neq i$, allowing for pairwise noise where similar classes are more frequently confused (e.g., a ”bird” class being misclassified as ”plane” rather than ”dog”).
NNAR noise is the most complex type, with mislabeling probabilities influenced by both instance features and true class. Here, the misclassification probability is defined as $p(\tilde{y}=j|y=i,x)$, where $x$ denotes instance features. In practice, NNAR noise simulates real-world scenarios where specific regions in feature space are more prone to errors, such as when instances are visually or statistically similar to those in other classes.
Figure 1 shows noise transition matrices for a dataset with four classes under no noise, 20% symmetric (NCAR), and 20% pair (NAR) noise. Given the complexity of generating NNAR noise, and for comparability with other studies, this work focuses on NCAR and NAR noise, hereafter referred to as ”symmetric” and ”pair” noise, respectively.